OpenStudy (anonymous):

Is this in simplest form? (x+2)(x-5) over (x-2)(5-x)

OpenStudy (solomonzelman):

\(\Large\color{orangered}{ \frac{(x+2)(x-5) }{(x-2)(5-x)} }\) \(\Large\color{orangered}{ \frac{(x+2)(x-5) }{(x-2)(-1)(x-5)} }\) \(\Large\color{orangered}{ \frac{x+2 }{(x-2)(-1)} }\) \(\Large\color{orangered}{ \frac{x+2 }{-x+2} }\)

OpenStudy (anonymous):

I thought that is what you did. I remember having the multiply by 1 to get the (5-x) into the right format. Thanks!

OpenStudy (solomonzelman):

yes, just multiplying by -1 and canceling the terms out :)